# Chain-linked Multiple Matrix Integration via Embedding Alignment

### v2

@@ D.8 Extension to symmetric indefinite matrices
$\|\mathring{\mathbf{W}}^{(i)\top}\mathring{\mathbf{W}}^{(i,j)}\mathring{\mathbf{W}}^{(j)}-\mathbf{Q}^{(i)}(\mathbf{Q}^{(j)})^{-1}\|$ $\leq\|\hat{\mathbf{Y}}^{\dagger}-\mathbf{Y}^{\dagger}\|\cdot\|\hat{\mathbf{Z}}-\mathbf{Z}\|+\|\mathbf{Y}^{\dagger}(\hat{\mathbf{Z}}-\mathbf{Z})\|$ (D.35)
$+\|\hat{\mathbf{Y}}^{\dagger}-\mathbf{Y}^{\dagger}\|\cdot\|\hat{\mathbf{Y}}-\mathbf{Y}\|\cdot\|\mathbf{Y}^{\dagger}\|\cdot\|\mathbf{Z}\|+\|\mathbf{Y}^{\dagger}(\hat{\mathbf{Y}}-\mathbf{Y})\|\cdot\|\mathbf{Y}^{\dagger}\|\cdot\|\mathbf{Z}\|.$
We now bound the terms on the right side of Eq. (D.35). By Lemma D.12 we have
$\|\hat{\mathbf{Y}}-\mathbf{Y}\|\lesssim n_{i,j}^{1/2}\cdot\|\hat{\mathbf{X}}^{(i)}\mathring{\mathbf{W}}^{(i)}-\mathbf{X}^{(i)}\|_{2\to\infty}\lesssim\frac{n_{i,j}^{1/2}\gamma_{i}}{(q_{i}n_{i}\mu_{i})^{1/2}},$ (D.36)
$\|\hat{\mathbf{Z}}-\mathbf{Z}\|\lesssim n_{i,j}^{1/2}\cdot\|\hat{\mathbf{X}}^{(j)}\mathring{\mathbf{W}}^{(j)}-\mathbf{X}^{(j)}\|_{2\to\infty}\lesssim\frac{n_{i,j}^{1/2}\gamma_{j}}{(q_{j}n_{j}\mu_{j})^{1/2}}$
@@
$\vartheta_{i,j}:=\|\mathbf{X}^{(j)}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}\|^{2}=\|\mathbf{Z}\|^{2},\quad\theta_{i,j}:=\lambda_{d}(\mathbf{X}^{(i)\top}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}\mathbf{X}^{(i)}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle})=\lambda_{d}(\mathbf{Y}^{\top}\mathbf{Y})$
Thus $\|{\mathbf{Y}}^{\dagger}\|=\theta_{i,j}^{-1/2}.$ Because $\|\hat{\mathbf{Y}}-\mathbf{Y}\|\ll\lambda_{d}^{1/2}(\mathbf{Y}^{\top}\mathbf{Y})$ with high probability, we also have
$\|\hat{\mathbf{Y}}^{\dagger}\|=\lambda_{d}^{-1/2}(\hat{\mathbf{Y}}^{\top}\hat{\mathbf{Y}})\leq 2\lambda_{d}^{-1/2}(\mathbf{Y}^{\top}\mathbf{Y})\lesssim\theta_{i,j}^{-1/2}$
@@
$\|\hat{\mathbf{Y}}^{\dagger}-\mathbf{Y}^{\dagger}\|\leq\sqrt{2}\|\hat{\mathbf{Y}}^{\dagger}\|\cdot\|\mathbf{Y}^{\dagger}\|\cdot\|\hat{\mathbf{Y}}-\mathbf{Y}\|\lesssim\frac{n_{i,j}^{1/2}\gamma_{i}}{\theta_{i,j}(q_{i}n_{i}\mu_{i})^{1/2}}$ (D.37)
with high probability. Next, by Eq. (D.32), we have
$\|\mathbf{Y}^{\dagger}(\hat{\mathbf{Z}}-\mathbf{Z})\|\leq\|(\mathbf{Y}^{\top}\mathbf{Y})^{-1}\|\cdot\|\mathbf{Y}^{\top}\mathbf{E}^{(i)}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}\mathbf{X}^{(i)}(\mathbf{X}^{(i)\top}\mathbf{X}^{(i)})^{-1}\|+\|\mathbf{Y}^{\dagger}\|\cdot\|\mathbf{R}^{(i)}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}\|,$
$\|\mathbf{Y}^{\dagger}(\hat{\mathbf{Y}}-\mathbf{Y})\|\leq\|(\mathbf{Y}^{\top}\mathbf{Y})^{-1}\|\cdot\|\mathbf{Y}^{\top}\mathbf{E}^{(j)}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}\mathbf{X}^{(j)}(\mathbf{X}^{(j)\top}\mathbf{X}^{(j)})^{-1}\|+\|\mathbf{Y}^{\dagger}\|\cdot\|\mathbf{R}^{(j)}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}\|.$
By similar results to Lemma D.10 we have
@@
$\|\mathbf{Y}^{\dagger}(\hat{\mathbf{Z}}-\mathbf{Z})\|\lesssim\frac{n_{i,j}^{1/2}\gamma_{i}\|\mathbf{X}^{(i)}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}\|_{2\to\infty}}{\theta_{i,j}(q_{i}n_{i}\mu_{i})^{1/2}}+\frac{n_{i,j}^{1/2}}{\theta_{i,j}^{1/2}}\Bigl{(}\frac{\gamma_{i}^{2}}{q_{i}n_{i}\mu_{i}^{3/2}}+\frac{\gamma_{i}}{q_{i}^{1/2}n_{i}\mu_{i}^{1/2}}\Bigr{)},$ (D.38)
$\|\mathbf{Y}^{\dagger}(\hat{\mathbf{Y}}-\mathbf{Y})\|\lesssim\frac{n_{i,j}^{1/2}\gamma_{j}\|\mathbf{X}^{(i)}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}\|_{2\to\infty}}{\theta_{i,j}(q_{j}n_{j}\mu_{j})^{1/2}}+\frac{n_{i,j}^{1/2}}{\theta_{i,j}^{1/2}}\Bigl{(}\frac{\gamma_{j}^{2}}{q_{j}n_{j}\mu_{j}^{3/2}}+\frac{\gamma_{j}}{q_{j}^{1/2}n_{j}\mu_{j}^{1/2}}\Bigr{)}$
with high probability. Combining Eq. (D.35), Eq. (D.36), Eq. (D.37), and Eq. (D.38) we have the desired error rate of $\|\mathring{\mathbf{W}}^{(i)\top}\mathring{\mathbf{W}}^{(i,j)}\mathring{\mathbf{W}}^{(j)}-\mathbf{Q}^{(i)}(\mathbf{Q}^{(j)})^{-1}\|$. ∎
We now prove Theorem 4. In the case of a chain $(i_{0},i_{1},\dots,i_{L})$ we have
$\mathring{\mathbf{W}}^{(i_{0})\top}\Bigl{(}\prod_{\ell=1}^{L}\mathring{\mathbf{W}}^{(i_{\ell-1},i_{\ell})}\Bigr{)}\mathring{\mathbf{W}}^{(i_{L})}=\prod_{\ell=1}^{L}\mathring{\mathbf{W}}^{(i_{\ell-1})\top}\mathring{\mathbf{W}}^{(i_{\ell-1},i_{\ell})}\mathring{\mathbf{W}}^{(i_{\ell})}=\prod_{\ell=1}^{L}\tilde{\mathbf{T}}^{(i_{\ell-1},i_{\ell})},$
where $\prod$ is matrix product and $\tilde{\mathbf{T}}^{(i_{\ell-1},i_{\ell})}:=\mathring{\mathbf{W}}^{(i_{\ell-1})\top}\mathring{\mathbf{W}}^{(i_{\ell-1},i_{\ell})}\mathring{\mathbf{W}}^{(i_{\ell})}$. Furthermore, for any $2\leq\ell\leq L$ we also have
@@
$\Bigl{(}\prod_{k=1}^{\ell}\tilde{\mathbf{T}}^{(i_{k-1},i_{k})}\Bigr{)}-\mathbf{Q}^{(i_{0})}(\mathbf{Q}^{(i_{\ell})})^{-1}$ $=\Bigl{(}\prod_{k=1}^{\ell-1}\tilde{\mathbf{T}}^{(i_{k-1},i_{k})}\Bigr{)}\tilde{\mathbf{T}}^{(i_{\ell-1},i_{\ell})}-\mathbf{Q}^{(i_{0})}(\mathbf{Q}^{(i_{\ell-1})})^{-1}\mathbf{Q}^{(i_{\ell}-1)}(\mathbf{Q}^{(i_{\ell})})^{-1}$ (D.39)
$=\Bigl{(}\Bigl{(}\prod_{k=1}^{\ell-1}\tilde{\mathbf{T}}^{(i_{k-1},i_{k})}\Bigr{)}-\mathbf{Q}^{(i_{0})}(\mathbf{Q}^{(i_{\ell-1})})^{-1}\Bigr{)}(\tilde{\mathbf{T}}^{(i_{\ell-1},i_{\ell})}-\mathbf{Q}^{(i_{\ell-1})}(\mathbf{Q}^{(i_{\ell})})^{-1})$
$+\Bigl{(}\Bigl{(}\prod_{k=1}^{\ell-1}\tilde{\mathbf{T}}^{(i_{k-1},i_{k})}\Bigr{)}-\mathbf{Q}^{(i_{0})}(\mathbf{Q}^{(i_{\ell-1})})^{-1}\Bigr{)}\mathbf{Q}^{(i_{\ell-1})}(\mathbf{Q}^{(i_{\ell})})^{-1}$
$+\mathbf{Q}^{(i_{0})}(\mathbf{Q}^{(i_{\ell-1})})^{-1}\big{(}\tilde{\mathbf{T}}^{(i_{\ell-1},i_{\ell})}-\mathbf{Q}^{(i_{\ell-1})}(\mathbf{Q}^{(i_{\ell})})^{-1}\big{)}.$
Define $a_{\ell}:=\|\prod_{\ell=1}^{L}\tilde{\mathbf{T}}^{(i_{\ell-1},i_{\ell})}-\mathbf{Q}^{(i_{0})}(\mathbf{Q}^{(i_{\ell})})^{-1}\|$ for $1\leq\ell\leq L$. We then have $a_{1}\leq\alpha_{i_{0},i_{1}}$ with high probability by Lemma D.13, and have
@@
for $2\leq\ell\leq L$ with high probability by Eq. (D.39), where
@@
$\begin{split}\hat{\mathbf{P}}_{\mathcal{U}_{i_{0}},\mathcal{U}_{i_{L}}}-\mathbf{P}_{\mathcal{U}_{i_{0}},\mathcal{U}_{i_{L}}}&=\hat{\mathbf{X}}^{(i_{0})}\bigl{(}\prod_{\ell=1}^{L}\mathring{\mathbf{W}}^{(i_{\ell-1},i_{\ell})}\bigr{)}\mathbf{I}_{d_{+},d_{-}}\hat{\mathbf{X}}^{(i_{L})\top}-\mathbf{X}_{\mathcal{U}_{i_{0}}}\mathbf{I}_{d_{+},d_{-}}\mathbf{X}_{\mathcal{U}_{i_{L}}}^{\top}\\ &=\hat{\mathbf{X}}^{(i_{0})}\mathring{\mathbf{W}}^{(i_{0})}\bigl{(}\prod_{\ell=1}^{L}\tilde{\mathbf{T}}^{(i_{\ell-1},i_{\ell})}\bigr{)}\mathring{\mathbf{W}}^{(i_{L})\top}\mathbf{I}_{d_{+},d_{-}}\hat{\mathbf{X}}^{(i_{L})\top}-\mathbf{X}^{(i_{0})}\mathbf{Q}^{(i_{0})}\mathbf{I}_{d_{+},d_{-}}\mathbf{Q}^{(i_{L})\top}\mathbf{X}^{(i_{L})\top}\\ &=\hat{\mathbf{X}}^{(i_{0})}\mathring{\mathbf{W}}^{(i_{0})}\bigl{(}\prod_{\ell=1}^{L}\tilde{\mathbf{T}}^{(i_{\ell-1},i_{\ell})}\bigr{)}\mathbf{I}_{d_{+},d_{-}}\mathring{\mathbf{W}}^{(i_{L})\top}\hat{\mathbf{X}}^{(i_{L})\top}-\mathbf{X}^{(i_{0})}\mathbf{Q}^{(i_{0})}(\mathbf{Q}^{(i_{L})})^{-1}\mathbf{I}_{d_{+},d_{-}}\mathbf{X}^{(i_{L})\top},\end{split}$
where the last equality follows from the facts that $\mathring{\mathbf{W}}^{(i_{L})}\in\mathcal{O}_{d}\cap\mathcal{O}_{d_{+},d_{-}}$ and $\mathbf{Q}^{(i_{L})}\in\mathcal{O}_{d_{+},d_{-}}$. Let $\xi_{i}=\hat{\mathbf{X}}^{(i)}\mathring{\mathbf{W}}^{(i)}-\mathbf{X}^{(i)}$ for $i\in\{i_{0},i_{L}\}$. Following the same derivations as that for Eq. (C.4), with Lemma D.12 replacing Lemma C.1, we obtain
$\begin{split}\hat{\mathbf{P}}_{\mathcal{U}_{i_{0}},\mathcal{U}_{i_{L}}}-\mathbf{P}_{\mathcal{U}_{i_{0}},\mathcal{U}_{i_{L}}}&=\mathbf{E}^{(i_{0})}\mathbf{X}^{(i_{0})}(\mathbf{X}^{(i_{0})\top}\mathbf{X}^{(i_{0})})^{-1}\mathbf{I}_{d_{+},d_{-}}\mathbf{Q}^{(i_{0})}(\mathbf{Q}^{(i_{L})})^{-1}\mathbf{I}_{d_{+},d_{-}}\mathbf{X}^{(i_{L})\top}\\ &+\mathbf{X}^{(i_{0})}\mathbf{Q}^{(i_{0})}(\mathbf{Q}^{(i_{L})})^{-1}(\mathbf{X}^{(i_{L})\top}\mathbf{X}^{(i_{L})})^{-1}\mathbf{X}^{(i_{L})\top}\mathbf{E}^{(i_{L})}\\ &+\mathbf{R}^{(i_{0},i_{L})}+\mathbf{S}^{(i_{0},i_{1},\dots,i_{L})},\end{split}$
@@
$\mathbf{S}^{(i_{0},i_{1},\dots,i_{L})}=(\mathbf{X}^{(i_{0})}+\xi_{i_{0}})(\prod_{\ell=1}^{L}\tilde{\mathbf{T}}^{(i_{\ell-1},i_{\ell})}-\mathbf{Q}^{(i_{0})}(\mathbf{Q}^{(i_{L})})^{-1})\mathbf{I}_{d_{+},d_{-}}(\mathbf{X}^{(i_{L})}+\xi_{i_{L}})^{\top}.$
Substituting the bounds for $\mathbf{R}^{(i)}$ and $\xi_{i}$ in Lemma D.12, we obtain

### v4

@@ E.8 Extension to symmetric indefinite matrices
$\|\mathring{\mathbf{W}}^{(i)\top}\mathring{\mathbf{W}}^{(i,j)}\mathring{\mathbf{W}}^{(j)}-\mathbf{Q}^{(i)}(\mathbf{Q}^{(j)})^{-1}\|$ $\leq\|\widehat{\mathbf{Y}}^{\dagger}-\mathbf{Y}^{\dagger}\|\cdot\|\widehat{\mathbf{Z}}-\mathbf{Z}\|+\|\mathbf{Y}^{\dagger}(\widehat{\mathbf{Z}}-\mathbf{Z})\|$ (E.35)
$+\|\widehat{\mathbf{Y}}^{\dagger}-\mathbf{Y}^{\dagger}\|\cdot\|\widehat{\mathbf{Y}}-\mathbf{Y}\|\cdot\|\mathbf{Y}^{\dagger}\|\cdot\|\mathbf{Z}\|+\|\mathbf{Y}^{\dagger}(\widehat{\mathbf{Y}}-\mathbf{Y})\|\cdot\|\mathbf{Y}^{\dagger}\|\cdot\|\mathbf{Z}\|.$
We now bound the terms on the right side of Eq. (E.35). By Lemma E.12 we have
$\|\widehat{\mathbf{Y}}-\mathbf{Y}\|\lesssim n_{i,j}^{1/2}\cdot\|\widehat{\mathbf{X}}^{(i)}\mathring{\mathbf{W}}^{(i)}-\mathbf{X}^{(i)}\|_{2\to\infty}\lesssim\frac{n_{i,j}^{1/2}\gamma_{i}}{(q_{i}n_{i}\mu_{i})^{1/2}},$ (E.36)
$\|\widehat{\mathbf{Z}}-\mathbf{Z}\|\lesssim n_{i,j}^{1/2}\cdot\|\widehat{\mathbf{X}}^{(j)}\mathring{\mathbf{W}}^{(j)}-\mathbf{X}^{(j)}\|_{2\to\infty}\lesssim\frac{n_{i,j}^{1/2}\gamma_{j}}{(q_{j}n_{j}\mu_{j})^{1/2}}$
@@
$\vartheta_{i,j}=\|\mathbf{X}^{(j)}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}\|^{2}=\|\mathbf{Z}\|^{2},\quad\theta_{i,j}=\lambda_{d}(\mathbf{X}^{(i)\top}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}\mathbf{X}^{(i)}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle})=\lambda_{d}(\mathbf{Y}^{\top}\mathbf{Y})$
Thus $\|{\mathbf{Y}}^{\dagger}\|=\theta_{i,j}^{-1/2}.$ Because $\|\widehat{\mathbf{Y}}-\mathbf{Y}\|\ll\lambda_{d}^{1/2}(\mathbf{Y}^{\top}\mathbf{Y})$ with high probability, we also have
$\|\widehat{\mathbf{Y}}^{\dagger}\|=\lambda_{d}^{-1/2}(\widehat{\mathbf{Y}}^{\top}\widehat{\mathbf{Y}})\leq 2\lambda_{d}^{-1/2}(\mathbf{Y}^{\top}\mathbf{Y})\lesssim\theta_{i,j}^{-1/2}$
@@
$\|\widehat{\mathbf{Y}}^{\dagger}-\mathbf{Y}^{\dagger}\|\leq\sqrt{2}\|\widehat{\mathbf{Y}}^{\dagger}\|\cdot\|\mathbf{Y}^{\dagger}\|\cdot\|\widehat{\mathbf{Y}}-\mathbf{Y}\|\lesssim\frac{n_{i,j}^{1/2}\gamma_{i}}{\theta_{i,j}(q_{i}n_{i}\mu_{i})^{1/2}}$ (E.37)
with high probability. Next, by Eq. (E.32), we have
$\|\mathbf{Y}^{\dagger}(\widehat{\mathbf{Z}}-\mathbf{Z})\|\leq\|(\mathbf{Y}^{\top}\mathbf{Y})^{-1}\|\cdot\|\mathbf{Y}^{\top}\mathbf{E}^{(i)}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}\mathbf{X}^{(i)}(\mathbf{X}^{(i)\top}\mathbf{X}^{(i)})^{-1}\|+\|\mathbf{Y}^{\dagger}\|\cdot\|\mathbf{R}^{(i)}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}\|$,
$\|\mathbf{Y}^{\dagger}(\widehat{\mathbf{Y}}-\mathbf{Y})\|\leq\|(\mathbf{Y}^{\top}\mathbf{Y})^{-1}\|\cdot\|\mathbf{Y}^{\top}\mathbf{E}^{(j)}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}\mathbf{X}^{(j)}(\mathbf{X}^{(j)\top}\mathbf{X}^{(j)})^{-1}\|+\|\mathbf{Y}^{\dagger}\|\cdot\|\mathbf{R}^{(j)}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}\|$.
By similar results to Lemma E.10 we have
@@
$\|\mathbf{Y}^{\dagger}(\widehat{\mathbf{Z}}-\mathbf{Z})\|\lesssim\frac{n_{i,j}^{1/2}\gamma_{i}\|\mathbf{X}^{(i)}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}\|_{2\to\infty}}{\theta_{i,j}(q_{i}n_{i}\mu_{i})^{1/2}}+\frac{n_{i,j}^{1/2}}{\theta_{i,j}^{1/2}}\Bigl{(}\frac{\gamma_{i}^{2}}{q_{i}n_{i}\mu_{i}^{3/2}}+\frac{\gamma_{i}}{q_{i}^{1/2}n_{i}\mu_{i}^{1/2}}\Bigr{)},$ (E.38)
$\|\mathbf{Y}^{\dagger}(\widehat{\mathbf{Y}}-\mathbf{Y})\|\lesssim\frac{n_{i,j}^{1/2}\gamma_{j}\|\mathbf{X}^{(i)}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}\|_{2\to\infty}}{\theta_{i,j}(q_{j}n_{j}\mu_{j})^{1/2}}+\frac{n_{i,j}^{1/2}}{\theta_{i,j}^{1/2}}\Bigl{(}\frac{\gamma_{j}^{2}}{q_{j}n_{j}\mu_{j}^{3/2}}+\frac{\gamma_{j}}{q_{j}^{1/2}n_{j}\mu_{j}^{1/2}}\Bigr{)}$
with high probability. Combining Eq. (E.35), Eq. (E.36), Eq. (E.37), and Eq. (E.38) we have the desired error rate of $\|\mathring{\mathbf{W}}^{(i)\top}\mathring{\mathbf{W}}^{(i,j)}\mathring{\mathbf{W}}^{(j)}-\mathbf{Q}^{(i)}(\mathbf{Q}^{(j)})^{-1}\|$. ∎
We now prove Theorem C.1. In the case of a chain $(i_{0},i_{1},\dots,i_{L})$ we have
$\mathring{\mathbf{W}}^{(i_{0})\top}\Bigl{(}\prod_{\ell=1}^{L}\mathring{\mathbf{W}}^{(i_{\ell-1},i_{\ell})}\Bigr{)}\mathring{\mathbf{W}}^{(i_{L})}=\prod_{\ell=1}^{L}\mathring{\mathbf{W}}^{(i_{\ell-1})\top}\mathring{\mathbf{W}}^{(i_{\ell-1},i_{\ell})}\mathring{\mathbf{W}}^{(i_{\ell})}=\prod_{\ell=1}^{L}\widetilde{\mathbf{T}}^{(i_{\ell-1},i_{\ell})},$
where $\prod$ is matrix product and $\widetilde{\mathbf{T}}^{(i_{\ell-1},i_{\ell})}:=\mathring{\mathbf{W}}^{(i_{\ell-1})\top}\mathring{\mathbf{W}}^{(i_{\ell-1},i_{\ell})}\mathring{\mathbf{W}}^{(i_{\ell})}$. Furthermore, for any $2\leq\ell\leq L$ we also have
@@
$\Bigl{(}\prod_{k=1}^{\ell}\widetilde{\mathbf{T}}^{(i_{k-1},i_{k})}\Bigr{)}-\mathbf{Q}^{(i_{0})}(\mathbf{Q}^{(i_{\ell})})^{-1}$ $=\Bigl{(}\prod_{k=1}^{\ell-1}\widetilde{\mathbf{T}}^{(i_{k-1},i_{k})}\Bigr{)}\widetilde{\mathbf{T}}^{(i_{\ell-1},i_{\ell})}-\mathbf{Q}^{(i_{0})}(\mathbf{Q}^{(i_{\ell-1})})^{-1}\mathbf{Q}^{(i_{\ell}-1)}(\mathbf{Q}^{(i_{\ell})})^{-1}$ (E.39)
$=\Bigl{(}\Bigl{(}\prod_{k=1}^{\ell-1}\widetilde{\mathbf{T}}^{(i_{k-1},i_{k})}\Bigr{)}-\mathbf{Q}^{(i_{0})}(\mathbf{Q}^{(i_{\ell-1})})^{-1}\Bigr{)}(\widetilde{\mathbf{T}}^{(i_{\ell-1},i_{\ell})}-\mathbf{Q}^{(i_{\ell-1})}(\mathbf{Q}^{(i_{\ell})})^{-1})$
$+\Bigl{(}\Bigl{(}\prod_{k=1}^{\ell-1}\widetilde{\mathbf{T}}^{(i_{k-1},i_{k})}\Bigr{)}-\mathbf{Q}^{(i_{0})}(\mathbf{Q}^{(i_{\ell-1})})^{-1}\Bigr{)}\mathbf{Q}^{(i_{\ell-1})}(\mathbf{Q}^{(i_{\ell})})^{-1}$
$+\mathbf{Q}^{(i_{0})}(\mathbf{Q}^{(i_{\ell-1})})^{-1}\big{(}\widetilde{\mathbf{T}}^{(i_{\ell-1},i_{\ell})}-\mathbf{Q}^{(i_{\ell-1})}(\mathbf{Q}^{(i_{\ell})})^{-1}\big{)}.$
Define $a_{\ell}:=\|\prod_{\ell=1}^{L}\widetilde{\mathbf{T}}^{(i_{\ell-1},i_{\ell})}-\mathbf{Q}^{(i_{0})}(\mathbf{Q}^{(i_{\ell})})^{-1}\|$ for $1\leq\ell\leq L$. We then have $a_{1}\leq\alpha_{i_{0},i_{1}}$ with high probability by Lemma E.13, and have
@@
for $2\leq\ell\leq L$ with high probability by Eq. (E.39), where
@@
$\begin{split}\widehat{\mathbf{P}}_{\mathcal{U}_{i_{0}},\mathcal{U}_{i_{L}}}-\mathbf{P}_{\mathcal{U}_{i_{0}},\mathcal{U}_{i_{L}}}&=\widehat{\mathbf{X}}^{(i_{0})}\bigl{(}\prod_{\ell=1}^{L}\mathring{\mathbf{W}}^{(i_{\ell-1},i_{\ell})}\bigr{)}\mathbf{I}_{d_{+},d_{-}}\widehat{\mathbf{X}}^{(i_{L})\top}-\mathbf{X}_{\mathcal{U}_{i_{0}}}\mathbf{I}_{d_{+},d_{-}}\mathbf{X}_{\mathcal{U}_{i_{L}}}^{\top}\\ &=\widehat{\mathbf{X}}^{(i_{0})}\mathring{\mathbf{W}}^{(i_{0})}\bigl{(}\prod_{\ell=1}^{L}\widetilde{\mathbf{T}}^{(i_{\ell-1},i_{\ell})}\bigr{)}\mathring{\mathbf{W}}^{(i_{L})\top}\mathbf{I}_{d_{+},d_{-}}\widehat{\mathbf{X}}^{(i_{L})\top}-\mathbf{X}^{(i_{0})}\mathbf{Q}^{(i_{0})}\mathbf{I}_{d_{+},d_{-}}\mathbf{Q}^{(i_{L})\top}\mathbf{X}^{(i_{L})\top}\\ &=\widehat{\mathbf{X}}^{(i_{0})}\mathring{\mathbf{W}}^{(i_{0})}\bigl{(}\prod_{\ell=1}^{L}\widetilde{\mathbf{T}}^{(i_{\ell-1},i_{\ell})}\bigr{)}\mathbf{I}_{d_{+},d_{-}}\mathring{\mathbf{W}}^{(i_{L})\top}\widehat{\mathbf{X}}^{(i_{L})\top}-\mathbf{X}^{(i_{0})}\mathbf{Q}^{(i_{0})}(\mathbf{Q}^{(i_{L})})^{-1}\mathbf{I}_{d_{+},d_{-}}\mathbf{X}^{(i_{L})\top},\end{split}$
where the last equality follows from the facts that $\mathring{\mathbf{W}}^{(i_{L})}\in\mathcal{O}_{d}\cap\mathcal{O}_{d_{+},d_{-}}$ and $\mathbf{Q}^{(i_{L})}\in\mathcal{O}_{d_{+},d_{-}}$. Let $\xi_{i}=\widehat{\mathbf{X}}^{(i)}\mathring{\mathbf{W}}^{(i)}-\mathbf{X}^{(i)}$ for $i\in\{i_{0},i_{L}\}$. Following the same derivations as that for Eq. (D.4), with Lemma E.12 replacing Lemma D.1, we obtain
$\begin{split}\widehat{\mathbf{P}}_{\mathcal{U}_{i_{0}},\mathcal{U}_{i_{L}}}-\mathbf{P}_{\mathcal{U}_{i_{0}},\mathcal{U}_{i_{L}}}&=\mathbf{E}^{(i_{0})}\mathbf{X}^{(i_{0})}(\mathbf{X}^{(i_{0})\top}\mathbf{X}^{(i_{0})})^{-1}\mathbf{I}_{d_{+},d_{-}}\mathbf{Q}^{(i_{0})}(\mathbf{Q}^{(i_{L})})^{-1}\mathbf{I}_{d_{+},d_{-}}\mathbf{X}^{(i_{L})\top}\\ &+\mathbf{X}^{(i_{0})}\mathbf{Q}^{(i_{0})}(\mathbf{Q}^{(i_{L})})^{-1}(\mathbf{X}^{(i_{L})\top}\mathbf{X}^{(i_{L})})^{-1}\mathbf{X}^{(i_{L})\top}\mathbf{E}^{(i_{L})}\\ &+\mathbf{R}^{(i_{0},i_{L})}+\mathbf{S}^{(i_{0},i_{1},\dots,i_{L})},\end{split}$
@@
$\mathbf{S}^{(i_{0},i_{1},\dots,i_{L})}=(\mathbf{X}^{(i_{0})}+\xi_{i_{0}})(\prod_{\ell=1}^{L}\widetilde{\mathbf{T}}^{(i_{\ell-1},i_{\ell})}-\mathbf{Q}^{(i_{0})}(\mathbf{Q}^{(i_{L})})^{-1})\mathbf{I}_{d_{+},d_{-}}(\mathbf{X}^{(i_{L})}+\xi_{i_{L}})^{\top}.$
Substituting the bounds for $\mathbf{R}^{(i)}$ and $\xi_{i}$ in Lemma E.12, we obtain